OpenStudy (anonymous):

Use the quadratic formula need to show work 2x^2+3x-4=0

OpenStudy (anonymous):

\[\frac{ -b+\sqrt{b^2-4ac} }{ 2a }or \frac{ -b-\sqrt{b^2-4ac} }{ 2a }\]

OpenStudy (daffy_duck):

From equation, a = 2, b = 3, and c = -4

OpenStudy (daffy_duck):

Plug the number into s.mustafa88's equation

OpenStudy (daffy_duck):

*numbers